OpenStudy (anonymous):

A cyclist exerts a force of 15.0 N as he rides a bike 251 m in 30.0 s. How much power does the cyclist develop?

OpenStudy (anonymous):

P=w/t P is power, w is work, t is time W=f x d W is work, f is force, d is distance

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

Somewhat, thank you ^^

OpenStudy (anonymous):

P=15*251/30=125.5W